The length of a rectangle is 4 times the width. The area 144in squared. find the length and width of the rectangle.
Found 2 solutions by checkley77, richwmiller:
Answer by checkley77(12569)

L=4W
LW=144
4W*W=144
4W^2=144
W^2=144/4
W^2=36
W=SQRT36
W=6 ANS.
L=4*6=24 ANS.
PROOF:
6*24=144
144=144

Answer by richwmiller(9135)

The length of a rectangle is 4 times the width. The area 144in squared. find the length and width of the rectangle.
let w = width
l=length
l=4w
so we have length= 4w
and width =w
area equals length times width
144 sq in =4w*w
144=4w^2
36=w^2
6=w
width=6 inches
length=4w=24 inches